OpenStudy (anonymous):

Match each fraction with its simplified form. Matching Questions Matching Answers \frac{25}{30} 1. \frac{3}{5} \frac{22}{33} 2. \frac{5}{8} \frac{39}{65} 3. \frac{5}{6} \frac{10}{16} 4. \frac{2}{3}

OpenStudy (anonymous):

@zkoqmosii

OpenStudy (anonymous):

I am horrible at fractions and I don't want to fail math :/

OpenStudy (anonymous):

Could u help me more after this?

OpenStudy (zkoqmosii):

25/30 = 5/6 because 25 /5 = 5, and 30 / 5 = 6. 22/33 = 2/3 because 22 reduces to 2, while 33 reduces to 3. 39 / 65 = 3/5 because 39 can be reduced to 39, while 65 can be reduced to 5. 10 / 16 = 5 / 8 because 10 can be simplified by dividing it by 2, while 16 is reduced to 8. You can't reduce the 8 anymore because you have to do the same to the top and the bottom.